# Duoduo CLIP: Efficient 3D Understanding with Multi-View Images

### v1

@@ 4 Experiments
To compare our Duoduo CLIP to prior work, we conduct experiments showing shape classification on a collection of 3D assets (on Objaverse-LVIS in Section 4.2), real-world multi-view images (on MVImgNet in Section 4.3), and more fine-grained text-to-shape retrieval (Section 4.4). We first describe the experimental setup including implementation details and the datasets we use.
@@ 4.1 Experimental setup
3D Datasets. We follow OpenShape [25] and train our model on a combination of four datasets consisting of Objaverse [7], ABO [6], ShapeNet [3] and 3D-FUTURE [12]. Text for the shapes consists of captions produced by image-captioning models and filtered text descriptions from metadata of the 3D objects crawled from the web (see Liu et al. [25] for details). In total, the combined dataset contains $\sim$874k shapes, with 46k shapes within the LVIS subset of Objaverse which are used for evaluation. As in OpenShape [25], we consider two variations of the data: 1) Ensembled which contains all shapes, and 2) Ensembled (no LVIS) which excludes shapes from the LVIS subset.
@@
Renderings. We use images from Zero123 [26], which provides renderings of most shapes from Objaverse from $12$ random views using spherical sampling. The shapes not included in Zero123 (a few remaining objects in Objaverse and the other 3 datasets) are rendered using their Blender  [^1] https://www.blender.org/ script to match the shapes in OpenShape. We render $12$ views for all these objects.
Real-world data. MVImgNet [53] consists of multi-view images for 220k real-world objects over 238 categories. We first filter objects that has at least $12$ multi-views to obtain $\sim$190k remaining objects, where $12$ views are evenly sampled. BLIP-2 [23] is then used to produce captions for each of the views. To compare with point cloud based methods, we use the MVPNet subset containing $\sim$87k objects with densely reconstructed point clouds. After the preprocessing and filtering above, we obtain from MVPNet, $\sim$66k objects for the train split and $\sim$16k objects for the validation set, which has 180 categories. Note that the objects in the training and validation splits for MVPNet are disjoint but are both included in MVImgNet.
@@
Table 2: Classification accuracy on Objaverse-LVIS trained on Ensembled (no LVIS) and Ensembled. The zero shot CLIP (B-32 and BigG-14) models are not trained on 3D data and serve as baselines. For each model, we indicate the representation used for computing the embeddings, multi-view (MV) or point-cloud (PC), and the shape encoder (Enc) architecture that is used. Our multi-view based Duoduo CLIP is able to outperform point-cloud based approaches with the exception of Uni3D is a large model with 1B parameters. $\dagger$ indicates our reimplementation of OpenShape which is faster and outperforms the original OpenShape.
Ensembled (no LVIS) Ensembled
Method Rep Enc Top1 Top3 Top5 Top1 Top3 Top5
ZS B-32 (12F) MV Avg 35.7 54.8 62.1 35.7 54.8 62.1
ZS BigG-14 (12F) MV Avg 43.5 64.2 71.3 43.5 64.2 71.3
ULIP [49] PC PointBERT 21.4 38.1 46.0 26.8 44.8 52.6
OpenShape [25] PC PointBERT 39.1 60.8 68.9 46.8 69.1 77.0
@@
TAMM [58] PC PointBERT 42.0 63.6 71.7 50.7 73.2 80.6
Uni3D [60] PC 3D VIT 47.2 68.8 76.1 55.3 76.7 82.9
ShapeLLM [33] PC RECON++ 53.7 75.8 82.0
VIT-LENS [22] PC VIT-LENSG 50.1 71.3 78.1 52.0 73.3 79.9
Duoduo CLIP (5F) MV MVA 51.3 73.1 79.9 54.1 76.0 82.4
Duoduo CLIP (12F) MV MVA 52.7 74.5 81.3 55.2 77.1 83.4
4.2 Classification on Objaverse-LVIS
@@
We compare the performance of our model against two CLIP models for different numbers of frames in Figure 3, and recent state-of-the-art models in Table 2. For the CLIP models, we selected VIT with B-32 (the model that our Duoduo CLIP is initialized with), and BigG-14 (the CLIP teacher model used by OpenShape). Both models are run in a zero-shot manner without training on additional 3D data (the weights are frozen, with embeddings from the image encoder averaged).
@@
How well does performance scale with number of views? Having a more complete view of a 3D object is important for 3D representation learning as elements of the object may be occluded for certain views. Thus we investigate how the number of views affects the classification accuracy. From Figure 3 (left), we see that our model performance scales better as the amount of information (views) increases compared to the zero-shot CLIP models ($11.22$ for our Duoduo CLIP, and $7.45$ for B-32, the model we used for initializing our weights). This suggests that our training strategy of randomly sampling 1 to 6 views during training is effective in letting the model learn to leverage the number of views given to it. Although, the accuracy increases as the number of multi-views increases, the inference time also increases (Figure 3 right). As the accuracy tapers off as we increase the number of views, to avoid additional computational cost during inference, we use 12 views as the default.
@@
How many multi-views to beat a point cloud? Compared with using multi-views, using a point cloud representation offers good coverage of the entire geometry of the 3D object. Thus, we ask the question of how many views does it take to match existing methods on the Objaverse-LVIS classification task. We compare against ULIP [49], OpenShape [25], TAMM [58] that use point clouds with PointBERT [54] as the 3D encoder, but different ways of fine-tuning the 3D encoder. We also compare against more recent methods [60, 33, 22] that are the current top-performing methods on Objaverse-LVIS. From Table 2, on the Ensembled dataset, we are able to beat most previous methods with 5 multi-views except for Uni3D which has a much larger model (1B parameters). When we increase to 12 frames we are on par and even surpass Uni3D. This is also while using much fewer resources as seen in Table 1.
@@
Does multi-views generalize better compared to point cloud? From the Ensembled (no LVIS) column of Table 2, we see that even the zero-shot BigG-14 model used as the teacher model for OpenShape beats it. Suggesting that the CLIP models already generalize well to unseen data, and point cloud methods potentially having worse generalization on unseen shapes. With our model that is further trained on Ensembled (no LVIS) we surpass point cloud methods by a large margin. Showing that fine-tuning the CLIP model better preserves their generalization ability.
Table 3: Multi-view attention (MVA) ablation of accuracy on Objaverse-LVIS.
Method Frames Top 1 Top 3 Top 5
-MVA 6 54.22 76.13 82.74
+MVA 6 54.59 76.46 82.89
-MVA 12 55.02 76.92 83.49
+MVA 12 55.32 77.08 83.49
Table 4: Number of layers ablation. Accuracy on Objaverse-LVIS.
Method Top 1 Top 3 Top 5
3 layers 53.77 75.8 82.41
6 layers 55.24 77.08 83.43
12 layers (full) 55.32 77.08 83.49
Does multi-view attention help? We verify the effectiveness of our cross view attention in Table 4. For both -MVA and +MVA, we train all 12 attention layers, with the difference being +MVA has the cross-views attention enabled. We see that using multi-view attention helps the performance most in boosting Top1 accuracy for both 6-frame and 12-frame evaluations.
Do we need to train all the layers? We conduct an ablation to see if it is necessary to train all 12 layers. In Table 4, we show the accuracy for restricting the training to the last 3 and 6 MVA layers vs the full 12 MVA layers. We see that the number of trainable layers can be cut by half with minimal losses to accuracy, while further cuts start to hurt model performance. Training only half the layers reduces the VRAM usage by roughly half and allows us to train on GPUs with less memory. By restricting the training to the last 6 MVA layers, we can shift from training on 4$\times$×A40s (48 GB) to 4$\times$×A5000s (24 GB) (see Table 1).
Table 5: MVImgNet classification comparison.
@@
Table 6: Text to shape retrieval comparison. Methods are trained on the Text2Shape [4] dataset (T2S) or Ensembled.
@@
Text2Shape [4] T2S 0.40 2.37
Y2Seq2Seq [14] T2S 2.93 9.23
Parts2Words [41] T2S 12.72 32.98
TriCoLo [36] T2S 12.22 32.23
MXM-CLR [45] T2S 16.83 39.06
@@ 4.3 Classification on real-world data
Classification results. We report accuracy results on MVPNet in Table 6. For OpenShape, the top-1 accuracy is extremely low (only $3.55\%$ top-1) when evaluating on MVPNet directly with the model trained on Ensembled only. We suspect this is due to noisy reconstructed point clouds and inconsistent orientations compared to the training data. After training on MVPNet data, OpenShape’s performance improves significantly. For our Duoduo CLIP, even without training on MVImgNet or MVPNet data, our model achieves a top-1 accuracy of $49.16\%$. However, this is lower than the zero-shot CLIP B-32 model that our model is initialized from, suggesting possible overfitting to synthetic data. When we include training with multi-view images from MVPNet, our method surpasses both OpenShape and the zero-shot CLIP B-32 model even when just using one 1 view. Performance is further boosted when trained on the entire MVImgNet, highlighting the advantages of our Duoduo CLIP which uses multi-view images over OpenShape’s point clouds. By using multi-view images, we have better generalization on unseen data due to the priors from initializing with CLIP B-32. We also can train on more data, as multi-view images are more readily available compared to point clouds.
@@
Retrieval examples. In Figure 4, we present example retrieval results for querying Objaverse objects using images from the MVPNet validation set. We compare retrievals using OpenShape (+MVPNet) and Duoduo CLIP (+MVPNet). For OpenShape, we use 12 query images for each object, encode them with the frozen BigG-14 CLIP model, and average the image embeddings to obtain the query embedding. [^2] For the largest BigG-14 CLIP model, OpenShape does not use any linear projection layer $g^{I}$. This embedding is then matched against a database of point cloud embeddings of Objaverse objects to retrieve the top 3 results. For our method, Duoduo CLIP encodes both the query and Objaverse objects using 12 multi-view images.
@@
From Figure 4, we see that for basic objects like benches, both methods are able to retrieve plausible objects. For the armchair (row 2), while OpenShape retrieved objects of the same semantic class, our Duoduo CLIP retrieved armchairs that has better matching textures. This is because multi-view images can represent textures better than point clouds. Note that some retrieved objects look similar due to duplicate objects in Objaverse. For the jacket (row 3), the objects returned by OpenShape have broken geometry and are incoherent. We hypothesize this may happen when the sampled points produce embeddings that are by chance similar to the query image. This issue is also present last example (a red and black soccer ball) where the first object retrieved by OpenShape has broken geometry. However, our method also fails in this case, as our model retrieves ladybugs with red wings and black dots that resemble the soccer ball. Interestingly, OpenShape is able to retrieve a soccer ball and a red ball. This can indicate that point cloud based methods may be more robust to geometric properties of the object compared to the rendered images. However, we find that overall the retrieved objects from our method is more consistent.

### v3

@@ 4 Experiments
To compare our Duoduo CLIP to prior work, we conduct experiments showing shape classification on a collection of 3D assets (on Objaverse-LVIS in Sec. 4.2.1), real-world multi-view images (on MVImgNet in Sec. 4.2.2), and more fine-grained text-to-shape retrieval (Sec. 4.3). We also show image-to-shape retrieval results for out-of-distribution images (Sec. A.4.1) and concept mixing (Sec. A.4.2, where we retrieve shapes that maximizes similarity to two input images. We first describe the experimental setup including implementation details and the datasets we use.
@@ 4.1 Experimental setup
3D Datasets. We follow OpenShape (Liu et al., 2023a) and train our model on a combination of four datasets consisting of Objaverse (Deitke et al., 2023), ABO (Collins et al., 2022), ShapeNet (Chang et al., 2015) and 3D-FUTURE (Fu et al., 2021). Text for the shapes consists of captions produced by image-captioning models and filtered text descriptions from metadata of the 3D objects crawled from the web (see Liu et al. (2023a) for details). In total, the combined dataset contains $\sim$874k shapes, with 46k shapes within the LVIS subset of Objaverse which are used for evaluation. As in OpenShape (Liu et al., 2023a), we consider two variations of the data: 1) Ensembled which contains all shapes, and 2) Ensembled (no LVIS) which excludes shapes from the LVIS subset.
@@
Renderings. We use images from Zero123 (Liu et al., 2023b), which provides renderings of most shapes from Objaverse from $12$ random views using spherical sampling. The shapes not included in Zero123 (a few remaining objects in Objaverse and the other 3 datasets) are rendered using their Blender script to match the shapes in OpenShape. We render $12$ views for all these objects. Additional details and effects of view selection can be found in Sec. A.1.2.
Real-world data. Multi-view images. We leverage MVImgNet (Yu et al., 2023), a dataset comprising multi-view images for 220k real-world objects across 238 categories. To focus on objects with sufficient views, we filter those with at least $12$ views, evenly sampled, resulting in $\sim$190k objects. For each view, we generate captions using BLIP-2 (Li et al., 2023). To facilitate comparisons with point cloud-based methods, we use the MVPNet subset, which includes $\sim$87k objects with densely reconstructed point clouds. After preprocessing and filtering, we retain $\sim$66k objects for the training split and $\sim$16k for the validation set, spanning 180 categories. Importantly, the training and validation objects are disjoint but originate from the broader MVImgNet dataset. Reconstructed 3D data. To further evaluate our approach, we conduct experiments using objects from real-world scans. For instance, ScanObjectNN (Uy et al., 2019) provides point clouds derived from objects in SceneNN (Hua et al., 2016) and ScanNet (Dai et al., 2017). We evaluate on the test set, which contains 583 shapes across 15 categories, rendering 12 multi-view images directly from the point clouds (see Sec. A.3.1 for details). Additionally, we use the ScanNet validation set, consisting of 3825 objects across 17 classes, to perform classification tasks. For this, we crop multi-view images from the raw captured RGB frames in the dataset (details in Sec. A.3.2).
Figure 5: Left. We compare the Top 1 accuracies on Objaverse-LVIS for our model (Duoduo CLIP) and two pretrained CLIP models (zero-shot) for different number of frames. B-32 is the model we initialize from and train with and BigG-14 is the model used by OpenShape. The dashed lines show the difference between using 1 and 12 frames. Right. We show the Top1, Top3 and Top5 accuracies for our model for 1 to 12 frames. Gray dashed lines show the inference time(s) for 1 iteration and batch size 200 at different number of frames.
4.2 Classification Experiments
4.2.1 Objaverse-LVIS
@@
Table 2: Classification accuracy on Objaverse-LVIS (O-LVIS) and ScanObjectNN trained on Ensembled (no LVIS) and Ensembled. As baselines, the zero-shot CLIP models (ZS B-32 and ZS BigG-14) are not trained on 3D data, whereas the fine-tuned CLIP model (FT B-32) is trained using a single image per object. For each model, we indicate the representation used for computing the embeddings, multi-view (MV) or point-cloud (PC), and the shape encoder (Enc) architecture that is used. Our multi-view based Duoduo CLIP is able to outperform point-cloud based approaches with the exception of Uni3D is a large model with 1B parameters. $\dagger$ indicates our reimplementation of OpenShape which is faster and outperforms the original OpenShape. F represents the number of frames/multi-views used for evaluation.
Pretrain Dataset Ensembled (no LVIS) Ensembled
O-LVIS O-LVIS ScanObjectNN
Method Rep Enc Top1 Top3 Top5 Top1 Top3 Top5 Top1 Top3 Top5
ZS B-32 (12F) MV Avg 35.7 54.8 62.1 35.7 54.8 62.1 53.9 73.5 81.2
ZS BigG-14 (12F) MV Avg 43.5 64.2 71.3 43.5 64.2 71.3 56.7 78.2 85.8
FT B-32 (12F) MV Avg 50.1 72.0 79.2 53.0 74.7 81.4 55.1 75.6 83.9
ULIP (Xue et al., 2023) PC PointBERT 21.4 38.1 46.0 26.8 44.8 52.6 51.6 72.5 82.3
OpenShape (Liu et al., 2023a) PC PointBERT 39.1 60.8 68.9 46.8 69.1 77.0 52.2 79.7 88.7
@@
TAMM (Zhang et al., 2024) PC PointBERT 42.0 63.6 71.7 50.7 73.2 80.6 55.7 80.7 88.9
MixCon3D Gao et al. (2024) PC + MV PointBERT 47.5 69.0 76.2 52.5 74.5 81.2 58.6 80.3 89.2
Uni3D (Zhou et al., 2024) PC 3D VIT 47.2 68.8 76.1 55.3 76.7 82.9 65.3 85.5 92.7
ShapeLLM (Qi et al., 2024) PC RECON++ 53.7 75.8 82.0 65.4 84.1 89.7
VIT-LENS (Lei et al., 2024) PC VIT-LENSG 50.1 71.3 78.1 52.0 73.3 79.9 60.1 81.0 90.3
Duoduo CLIP (5F) MV MVA 51.3 73.1 79.9 54.1 76.0 82.4 60.7 82.4 88.5
Duoduo CLIP (12F) MV MVA 52.7 74.5 81.3 55.2 77.1 83.4 66.3 85.5 90.2
@@
We compare the performance of our model against two CLIP models for different numbers of frames in Fig. 5, and recent state-of-the-art models in Tab. 2. For the CLIP models, we selected VIT with B-32 (the model that our Duoduo CLIP is initialized with), and BigG-14 (the CLIP teacher model used by OpenShape). Both models are run in a zero-shot manner without training on additional 3D data (the weights are frozen, with embeddings from the image encoder averaged). We also include an additional baseline that mirrors our default model’s settings but is trained by sampling a single view per object. This setup simulates the single-image training of the original CLIP model, and we refer to it as FT B-32 in the tables.
@@
How well does performance scale with number of views? Having a more complete view of a 3D object is important for 3D representation learning as parts of the object may be occluded for some views. Thus we investigate how the number of views affects the classification accuracy. From Fig. 5 (left), we see that our model performance scales better as the amount of information (views) increases compared to the zero-shot CLIP models ($11.22$ for our Duoduo CLIP, and $7.45$ for B-32, the model we used for initializing our weights). This suggests that our training strategy of randomly sampling 1 to 6 views during training is effective in letting the model learn to leverage the number of views given to it. Although, the accuracy increases as the number of multi-views increases, the inference time also increases (Fig. 5 right). As the accuracy tapers off as we increase the number of views, to avoid additional computational cost during inference, we use 12 views as the default.
@@
How many multi-views to beat a point cloud? Compared with using multi-views, using a point cloud representation offers good coverage of the entire geometry of the 3D object. Thus, we ask the question of how many views does it take to match existing methods on the Objaverse-LVIS classification task. We compare against ULIP (Xue et al., 2023), OpenShape (Liu et al., 2023a), TAMM (Zhang et al., 2024) that use point clouds with PointBERT (Yu et al., 2022b) as the 3D encoder, but different ways of fine-tuning the 3D encoder. We also compare against more recent methods (Zhou et al., 2024, Qi et al., 2024, Lei et al., 2024) that are the current top-performing methods on Objaverse-LVIS. From Tab. 2, on the Ensembled dataset, we are able to beat most previous methods with 5 multi-views except for Uni3D which has a much larger model (1B parameters). When we increase to 12 frames we are on par and even surpass Uni3D. This is also while using much fewer resources as seen in Tab. 1.
@@
Does multi-views generalize better compared to point cloud? From the Ensembled (no LVIS) column of Tab. 2, we see that even the zero-shot BigG-14 model used as the teacher model for OpenShape beats it. Suggesting that the CLIP models already generalize well to unseen data, and point cloud methods potentially having worse generalization on unseen shapes. With our model that is further trained on Ensembled (no LVIS) we surpass point cloud methods by a large margin. Showing that fine-tuning the CLIP model better preserves their generalization ability. Additionally, our model generalizes well on ScanObjectNN (Uy et al., 2019) and outperforms point cloud methods on both Top1 and Top3 accuracies (see Sec. A.3 for more results on ScanObjectNN and ScanNet).
Table 3: MVPNet classification comparison.
@@
Table 4: Text to shape retrieval comparison. Methods are trained on the Text2Shape Chen et al. (2019) dataset (T2S) or Ensembled.
@@
Text2Shape Chen et al. (2019) T2S 0.40 2.37
Y2Seq2Seq Han et al. (2019) T2S 2.93 9.23
Parts2Words Tang et al. (2023) T2S 12.72 32.98
TriCoLo Ruan et al. (2024) T2S 12.22 32.23
MXM-CLR Wang et al. (2023) T2S 16.83 39.06
@@ 4.2.2 MVImgNet
Classification results. We report accuracy results on MVPNet in Tab. 4. For OpenShape, the top-1 accuracy is very low (only $10.8\%$ top-1) when evaluating on MVPNet directly with the model trained on Ensembled only. We suspect this is due to noisy reconstructed point clouds and inconsistent orientations compared to the training data. Note we re-align MVPNet point cloud’s up direction with Objaverse. After training on MVPNet data, OpenShape’s performance improves significantly. For our Duoduo CLIP, even without training on MVImgNet or MVPNet data, our model achieves a top-1 accuracy of $49.16\%$. However, this is lower than the zero-shot CLIP B-32 model that our model is initialized from, suggesting possible overfitting to synthetic data. When we include training with multi-view images from MVPNet, our method surpasses both OpenShape and the zero-shot CLIP B-32 model even when just using one 1 view. Performance is further boosted when trained on the entire MVImgNet, highlighting the advantages of our Duoduo CLIP which uses multi-view images over OpenShape’s point clouds. By using multi-view images, we have better generalization on unseen data due to the priors from initializing with CLIP B-32. We also can train on more data, as multi-view images are more readily available compared to point clouds.
@@
Retrieval examples. In Fig. 6, we present example retrieval results for querying Objaverse objects using images from the MVPNet validation set. We compare retrievals using OpenShape (+MVPNet) and Duoduo CLIP (+MVPNet). For OpenShape, we use 12 query images for each object, encode them with the frozen BigG-14 CLIP model, and average the image embeddings to obtain the query embedding. [^1] For the largest BigG-14 CLIP model, OpenShape does not use any linear projection layer $g^{I}$. This embedding is then matched against a database of point cloud embeddings of Objaverse objects to retrieve the top 3 results. For our method, Duoduo CLIP encodes both the query and Objaverse objects using 12 multi-view images.
@@
From Fig. 6, we see that for basic objects like benches, both methods are able to retrieve plausible objects. For the armchair (row 2), while OpenShape retrieved objects of the same semantic class, our Duoduo CLIP retrieved armchairs that has better matching textures. This is because multi-view images can represent textures better than point clouds. Note that some retrieved objects look similar due to duplicate objects in Objaverse. For the jacket (row 3), the objects returned by OpenShape have broken geometry and are incoherent. We hypothesize this may happen when the sampled points produce embeddings that are by chance similar to the query image. This issue is also present last example (a red and black soccer ball) where the first object retrieved by OpenShape has broken geometry. However, our method also fails in this case, as our model retrieves ladybugs with red wings and black dots that resemble the soccer ball. Interestingly, OpenShape is able to retrieve a soccer ball and a red ball. This can indicate that point cloud based methods may be more robust to geometric properties of the object compared to the rendered images. However, we find that overall the retrieved objects from our method are more consistent.
4.2.3 Model Ablations
Table 5: Ablation on the number of layers for accuracy on Objaverse-LVIS using 12 input frames. Default model highlighted in gray.
Method Top 1 Top 3 Top 5
3 layers 53.77 75.8 82.41
6 layers 55.24 77.08 83.43
12 layers (full) 55.32 77.08 83.49
Table 6: Multi-view attention (MVA) ablation of accuracy on Objaverse-LVIS (O-LVIS), MVPNet and ScanObjectNN with 12 frames. Default model highlighted in gray.
Method Layers O-LVIS MVPNet ScanObjectNN
-MVA 6 54.61 47.87 58.77
+MVA 6 55.24 49.16 66.32
-MVA 12 55.02 43.75 56.83
+MVA 12 55.32 44.42 64.15
Do we need to train all the layers? We conduct an ablation to see if it is necessary to train all 12 layers. In Tab. 6, we show the accuracy for restricting the training to the last 3 and 6 MVA layers vs the full 12 MVA layers. We see that the number of trainable layers can be cut by half with minimal losses to accuracy, while further cuts start to hurt model performance. Training only half the layers reduces the VRAM usage by roughly half and allows us to train on GPUs with less memory. By restricting the training to the last 6 MVA layers, we can shift from training on 4$\times$×A40s (48 GB) to 4$\times$×A5000s (24 GB) (see Tab. 1). Also we find that training less layers also helps preserve the priors of the model and make generalization on other datasets better as seen in Tab. 6.
Does multi-view attention help? We verify the effectiveness of our cross view attention in Tab. 6. +MVA and -MVA denote whether the cross-views attention is enabled. On Objaverse-LVIS, while the +MVA mechanisms help less +0.3 when all 12 layers are trainable, it makes a larger impact when only training 6 layers +0.63. Making the +MVA mechanism an effective way to maintain performance while being more efficient compared to training all 12 layers. In addition, we find that the +MVA also helps in making the model generalize better on unseen datasets like MVPNet and ScanObjectNN compared to -MVA. Inference speeds for the models can be found in Sec. A.1.1.